Circular No. 7850

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

CBAT@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 E1 (SOHO)

Further to IAUC 7842, a Kreutz sungrazing comet has been found by S. Hönig on SOHO website C3 images (cf. MPEC 2002-E18).

2002 UT  α_{2000}  δ_{2000}
Mar. 2.096  23^{h}14.^{m}4  − 8^{o}26′

COMET 2002 E2

Independent visual discovery reports of a comet have been received from Douglas Snyder and Shigeki Murakami. Available observations:

2002 UT  α_{2000}  δ_{2000}  m_{1} Observer
Mar. 11.514  18^{h}55.^{m}2  −0^{o}46′  13.0 Snyder
11.806  18 55.5  −0 35  11  Murakami
12.12641 18 55 50.^{s}25 −0 16 18′′.1  Pravec
12.12834 18 55 50.39 −0 16 11.4  _{”}
12.13035 18 55 50.52 −0 16 04.1 12  _{”}
12.13812 18 55 51.01 −0 15 38.4  Tichý
12.13942 18 55 51.01 −0 15 32.9  _{”}
12.14541 18 55 51.55 −0 15 10.2  _{”}
12.14599 18 55 51.64 −0 15 07.7  _{”}
12.14677 18 55 51.63 −0 15 05.6  _{”}
12.49212 18 56 14.65 +0 05 21.0 12.8 Hale
12.49755 18 56 15.06 +0 05 40.0 12.2  _{”}
12.50569 18 56 15.54 +0 06 08.9 13.0  _{”}
12.51385 18 56 16.06 +0 06 36.5 12.2  _{”}
12.51778 18 56 16.43 +0 06 50.8 11.5  _{”}

D. Snyder (Palominas, AZ). 0.50-m f/5 reflector. Coma diameter ∼ 4′, diffuse with little central condensation. Motion ∼ 5′ toward northwest during Mar. 11.469–11.514 UT. Observed again Mar. 12.46–12.50.

S. Murakami (Matsunoyama, Niigata, Japan). 0.46-m f/4.5 reflector (68×).

Coma diameter 3′. Slight motion toward northeast in 30 min. P. Pravec and P. Kušnirák (Ondřejov). 0.65-m f/3.6 reflector + CCD. Coma

diameter 1′.1, well condensed; dense starfield. M. Tichý and J. Tichá (Kleť). 0.57-m f/5.2 reflector + CCD. Coma diam-

eter > 30′′. A. Hale (Cloudcroft, NM). 0.20-m reflector + CCD. With a 0.41-m reflector

visually at Mar. 12.46, m_{1} = 10.4, coma diameter 3′.

2002 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green